OpenStudy (anonymous):

Amanda, Trey, and Jose have a total of $117 in their wallets. Trey has 4 times what Jose has. Jose has 9 less than Amanda. How much do they have in their wallets

OpenStudy (anonymous):

A+T+J=117 T=4J J+9=A So, (J+9)+4J+J=117 6J=117-9=108 J=18, T=72, A=27